OpenStudy (anonymous):

prime factorization of 260

OpenStudy (anonymous):

2^2 * 5 * 13

OpenStudy (anonymous):

Tomas.A you're wrong. 130/2 = 65 not 75

OpenStudy (slaaibak):

260 = 26 * 10 26 = 13 * 2 10 = 5 * 2 therefore it is: \[13 * 2^{2} * 5\]

OpenStudy (anonymous):

260/2=130 130/2=75 65/5=13 13/13=1 2*2*5*13=260

OpenStudy (asnaseer):

I would suggest you use wolram alpha for questions like these. see: http://www2.wolframalpha.com/input/?i=260

OpenStudy (anonymous):

factor tree :)

OpenStudy (anonymous):

i know how to do i just didnt feel like doing it

OpenStudy (anonymous):